OpenStudy (anonymous):

help plz!! The prism below is made of cubes which measure 1/4 of a foot on one side. What is the volume?

OpenStudy (anonymous):

https://app108.studyisland.com/pics/160598vol224.png

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (welshfella):

how many cubes are there?

OpenStudy (studygurl14):

Count how many cubes make up the length, width, and height of teh prism. Tell me what you get?

OpenStudy (anonymous):

there is 16 cubes

OpenStudy (welshfella):

and what is the volume of each cube ( each side = 1/4 foot)

OpenStudy (anonymous):

idk how to find volume

OpenStudy (welshfella):

remember volume = s * s* s where s = length of 1 side

OpenStudy (welshfella):

here s = 1/4

OpenStudy (anonymous):

so 1/4*1/4*1/4

OpenStudy (anonymous):

@welshfella

OpenStudy (abbiegalie):

To find volume of a rectangular prism you would do width*length*height. (volume of a cube is s*s*s but we aren't trying to find volume of the cubes, but rather the prism) However before you can plug in something to "s" you need to calculate the side length. If each cube is 1/4 (and I will just say 0.25 ft), since its easier to deal with), then to find the width, length and height, you need to multiply 0.25 ft by the number of cubes along the side. The width (basically the depth of the prism), is 0.25 * 2, since there are 2 cubes. So 0.25 * 2 = 0.50 ft. Therefore the length is 0.50 ft. The height will be 0.50 ft also since it is made up of 2 cubes. The length is going to be 0.25*4 (there are 4 cubes). 0.25*4 = 1 ft. Okay so there you have it your length, width and height! To find the volume of this rectangular prism multiply 0.50 * 0.50 * 1.